Gerald is considering getting his master’s degree in business administration (MBA). He wants to be sure this degree will positively impact his career and earnings though. He randomly samples managers and executives at 10 companies to see if individuals with a MBA (n = 17, = $103,000, SD = 17,000) make more per year than other individuals (n = 23, = $92,000, SD = 12,000). What is the upper limit for the confidence interval for the difference obtained?
Answer #1

Ans:

pooled standard deviation=SQRT(((17-1)*17000^2+(23-1)*12000^2)/(17+23-2))=14319.66

standard error for difference=14319.66*SQRT((1/17)+(1/23))=4580.093

df=17+23-2=38

t*=tinv(0.05,38)=2.024

Margin of error=2.024*4580.093=9270.1

Point estimate=103000-92000=11000

Upper limit=11000+9270.1=20270.1